# MODEM: A Morton-Order Degradation Estimation Mechanism for Adverse Weather Image Recovery

### v1

@@ 5.1 Comparisons with State-of-the-arts
Qualitative comparison. We provide qualitative comparisons across diverse scenarios in Figs. 8 and 6. For image desnowing on Snow100K [35], MODEM effectively removes heavy snowflakes and visual artifacts that other models [78, 44, 54] struggle to address. For joint deraining and dehazing on Outdoor-Rain [29, 30], MODEM excels in restoring richer texture details and produces images with noticeably fewer artifacts. For raindrop removal on the Raindrop [46], MODEM again demonstrates superior detail preservation and artifact reduction. Furthermore, on real-world snowy images [35], MODEM achieves superior results as shown in Fig. 6 even without any additional fine-tuning, showcasing excellent generalization and real-world applicability, which can be attributed to its profound understanding and adaptive modeling of degradation characteristics. More cases can be found in Section A.5.
5.2 Ablation Studies
We evaluate the impact of the Morton-Order scan, DDEM, DAFM and DSAM. Morton-Order scan facilitates structured spatial feature processing within the SSM, enabling the model to better capture contextual dependencies and preserve local structural coherence. As shown in Table 5, configurations incorporating the Morton scan generally yield improved performance, underscoring its benefit in organizing spatial information for sequential modeling. DDEM provides the degradation priors $Z_{0}$ and $Z_{1}$ that guide the DAFM and DSAM. As indicated in Table 5, configurations lacking DDEM exhibit a significant performance drop. DAFM, which utilizes the global degradation prior $Z_{0}$, plays a crucial role in adaptively modulating features based on the overall estimated weather type and severity. DSAM, guided by the spatially adaptive kernel $Z_{1}$, allows the MOS2D to selectively focus on and modulate features pertinent to local degradation characteristics or specific image regions. While removing DSAM shows marginal gains on certain snow metrics, it impairs performance on tasks like raindrop removal, which demands strong local adaptation due to the highly local nature of raindrops, and on mixed rain&haze scenarios where intricate local texture recovery is challenging. Thus, DSAM, guided by the spatially adaptive kernel $Z_{1}$, is crucial for achieving robust, balanced performance across diverse weather conditions by guiding the network to selectively address these local characteristics. The combination of all components employed in full MODEM, consistently yields the best results. This result is further supported by t-SNE [57] in Fig. 7. Compared to Histoformer [54], MODEM exhibits significantly improved clustering of features corresponding to different weather types. This indicates that the combined action of the components enables MODEM to learn more discriminative and well-separated feature representations, which directly contributes to enhance restoration capabilities.
@@ A.1 Implement Details
MODEM is implemented in PyTorch [45] and trained on 4 NVIDIA RTX 3090 GPUs. The main backbone contains residual groups with [4, 4, 6, 8, 6, 4, 4] MDSLs sequentially, and a final refinement group with 4 MDSLs. Downsampling and upsampling are performed using PixelUnshuffle and PixelShuffle, respectively. The initial channel size is 36. The DDEM employs residual groups with 2 MDSLs each, operating with 96 channels. We train MODEM in two stages. In Stage 1, DDEM takes the channel-concatenated degraded image $I_{\text{LQ}}$ and ground-truth $I_{\text{GT}}$ as input. We use the AdamW optimizer [37] with a learning rate of $3\times 10^{-4}$, a weight decay of $1\times 10^{-4}$, and betas set to $(0.9,0.999)$. The learning rate is scheduled by a Cosine Annealing Restart Cyclic LR [36] scheduler with periods of [92k, 208k] iterations and restart weights of [1, 1]. The minimum learning rates, $\eta_{\text{min}}$, for these cycles are [$3\times 10^{-4},1\times 10^{-6}$]. Progressive training is adopted with patch sizes [128, 160, 256, 320, 384] and corresponding per-GPU mini-batch sizes [6, 4, 2, 1, 1] for [92k, 84k, 56k, 36k, 32k] iterations each, totaling 300k iterations. In Stage 2, DDEM processes only $I_{\text{LQ}}$, initializing parameters from Stage 1. The same AdamW optimizer settings and Cosine Annealing Restart Cyclic LR scheduler parameters are employed. Progressive training uses patch sizes [128, 160, 256, 320, 376] with per-GPU mini-batch sizes [8, 5, 2, 1, 1].
@@ A.3 Complexity Analysis
We report the parameters and inference time. The inference time is performed on a single Nvidia RTX 3090, with single 256 $\times$ 256 input image, detailed in Table 6.
@@
Methods WGWSNet WeatherDiff Histoformer MODEM
Time (ms) 24.83 1.67×$\times 10^{6}$ 109.07 92.86
@@
A.4 Limitations
While our proposed MODEM demonstrates state-of-the-art performance across a variety of adverse weather conditions, we acknowledge certain limitations. As illustrated in Fig. 9 of challenging real-world snow scenarios, MODEM, like other contemporary methods, can encounter difficulties in achieving perfect restoration when faced with images containing extremely large, high-contrast snowflakes. Such scenarios are particularly challenging if these specific visual patterns of snow, differing significantly in scale, density, or opacity from typical training examples, are underrepresented or entirely absent in the training data distribution. Nevertheless, empowered by its robust degradation estimation capabilities, MODEM still achieves comparatively better results in these extreme cases, effectively suppressing artifacts and preserving some structural detail. This highlights an ongoing challenge in achieving perfect generalization to all unseen severe degradations, but also underscores the significant benefit of incorporating explicit and adaptive degradation estimation.
A.5 More Visual Comparisons
Further visual comparisons are presented in Figs. 10, 11, 12 and 13.

### v2

@@ 6 Conclusion
Acknowledgement
We would like to thank Mingjia Li for the insightful discussions and feedback. This work was partially supported by the National Natural Science Foundation of China under Grant nos 62372251. The computational resources of this work was partially supported by TPU Research Cloud (TRC).
@@ A.1 Implement Details
MODEM is implemented in PyTorch [56] and trained on 4 NVIDIA RTX 3090 GPUs. The main backbone contains residual groups with [4, 4, 6, 8, 6, 4, 4] MDSLs sequentially, and a final refinement group with 4 MDSLs. Downsampling and upsampling are performed using PixelUnshuffle and PixelShuffle, respectively. The initial channel size is 36. The DDEM employs residual groups with 2 MDSLs each, operating with 96 channels. We train MODEM in two stages. In Stage 1, DDEM takes the channel-concatenated degraded image $I_{\text{LQ}}$ and ground-truth $I_{\text{GT}}$ as input. We use the AdamW optimizer [47] with a learning rate of $3\times 10^{-4}$, a weight decay of $1\times 10^{-4}$, and betas set to $(0.9,0.999)$. The learning rate is scheduled by a Cosine Annealing Restart Cyclic LR [46] scheduler with periods of [92k, 208k] iterations and restart weights of [1, 1]. The minimum learning rates, $\eta_{\text{min}}$, for these cycles are $3\times 10^{-4},1\times 10^{-6}$]. Progressive training is adopted with patch sizes [128, 160, 256, 320, 384] and corresponding per-GPU mini-batch sizes [6, 4, 2, 1, 1] for [92k, 84k, 56k, 36k, 32k] iterations each, totaling 300k iterations. In Stage 2, DDEM processes only $I_{\text{LQ}}$, initializing parameters from Stage 1. The same AdamW optimizer settings and Cosine Annealing Restart Cyclic LR scheduler parameters are employed. Progressive training uses patch sizes [128, 160, 256, 320, 376] with per-GPU mini-batch sizes [8, 5, 2, 1, 1].
@@ A.3 Complexity Analysis
We report the parameters and inference time. The inference time is performed on a single Nvidia RTX 3090, with single 256 $\times$ 256 input image, detailed in Table˜8.
@@
Methods WGWSNet [95] WeatherDiff [55] Histoformer [65] MODEM (Ours)
[…]
@@
Table 9: Comparison of inference time (ms) for different input sizes and average PSNR.
Input Size WGWSNet [95] WeatherDiff [55] Histoformer [65] MODEM (Ours)
[…]
Average PSNR 31.54 31.57 33.68 34.18
To be more comprehensive, we evaluate the inference speed on larger resolutions in Table˜9. As can be seen, compared to Transformer-based architectures like Histoformer [65], as the number of pixels increases, MODEM’s complexity scales in a near-linear fashion, demonstrating significantly better scalability. This stands in contrast to the quadratic complexity often associated with Transformers, granting MODEM a distinct computational advantage.
Table 10: Comparison of parameters and inference time for other Mamba-style methods.
Method MambaIR [20] FreqMamba [91] MODEM (Ours)
Parameters (M) 15.78 8.93 19.96
Time (ms) 790.61 233.41 92.86
Table 11: Comparison of inference time (ms) between Hilbert scan and Morton scan.
Input Size 256 $\times$ 256 512 $\times$ 512 1024 $\times$ 1024
Hilbert scan [28] 604.00 10134.91 88288.46
Morton scan (Ours) 92.86 443.02 1946.34
To further contextualize our model’s performance, we compared MODEM with other Mamba-style methods, MambaIR [20] and FreqMamba [91]. As shown in Table˜10, although MODEM has a larger parameter count compared to both MambaIR [20] and FreqMamba [91], it achieves a significantly faster inference time, highlighting the superior efficiency of our architectural design.
For scanning methods, Hilbert scan [28] is another locality-preserving alternative. However, it comes with a significantly higher computational cost, whereas the Morton-order can be calculated efficiently with simple bitwise operations as shown in our Eq.˜6. To be clear, we compared the inference speed of the Morton and Hilbert scans. For the Hilbert scan, we used the official implementation from LC-Mamba [28]. The speed (ms) of different resolutions are shown in the Table˜11, with all tests performed on a single RTX 3090 GPU. Our method is significantly faster than Hilbert.
A.4 More Visualizations of Features
A.5 Limitations
While our proposed MODEM demonstrates state-of-the-art performance across a variety of adverse weather conditions, we acknowledge certain limitations. As illustrated in Fig.˜10 of challenging real-world snow scenarios, MODEM, like other contemporary methods, can encounter difficulties in achieving perfect restoration when faced with images containing extremely large, high-contrast snowflakes. Such scenarios are particularly challenging if these specific visual patterns of snow, differing significantly in scale, density, or opacity from typical training examples, are underrepresented or entirely absent in the training data distribution. Nevertheless, empowered by its robust degradation estimation capabilities, MODEM still achieves comparatively better results in these extreme cases, effectively suppressing artifacts and preserving some structural detail. This highlights an ongoing challenge in achieving perfect generalization to all unseen severe degradations, but also underscores the significant benefit of incorporating explicit and adaptive degradation estimation.
@@ A.5 Limitations
A.6 More Ablation Studies
The Morton-order scans an image by processing it in small, contiguous block-like regions, moving from one adjacent block to the next (refer to the scale-space theory). This ensures that pixels that are close in the 2D image stay close in the 1D sequence. To further investigate the impact of the scanning scheme itself, we offer an ablation study within our MODEM, comparing the Morton scan against others [20, 74, 63, 25, 28, 24]. As shown in Table˜12, our Morton achieves the best performance.
Table 12: Ablation study of diffrent scanning scheme.
Methods Snow100K-L Snow100K-S Outdoor RainDrop
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
Raster (MambaIR [20]) 32.33 0.9283 38.03 0.9672 32.89 0.9399 32.69 0.9423
Continuous (Zigzag [74]) 32.17 0.9266 37.74 0.9662 32.35 0.9385 32.59 0.9422
OSS (VmambaIR [63]) 32.14 0.9262 37.39 0.9651 32.61 0.9387 32.11 0.9413
Local (LocalMamba [25]) 32.23 0.9266 37.75 0.9661 32.60 0.9382 32.53 0.9418
Hilbert (LC-Mamba [28]) 32.46 0.9287 37.96 0.9671 32.99 0.9414 32.82 0.9433
Morton (Ours) 32.52 0.9292 38.08 0.9673 33.10 0.9410 33.01 0.9434
We conduct an ablation study on the contribution of the loss terms. We have also performed an ablation study on the placement (before/after the average pooling) of the KL divergence loss. The results are presented in the table below. They confirm that each component contributes positively.
Table 13: Ablation study of the loss function.
Methods Snow100K-L Snow100K-S Outdoor RainDrop
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
w/o Correlation Loss 32.30 0.9278 37.79 0.9668 32.91 0.9403 32.69 0.9427
w/o KL Loss 32.12 0.9249 37.61 0.9650 32.37 0.9224 32.38 0.9390
Replace KL Loss 31.85 0.9237 37.16 0.9638 32.77 0.9389 32.57 0.9387
MODEM 32.52 0.9292 38.08 0.9673 33.10 0.9410 33.01 0.9434
A.7 More Quantitative Comparisons
We retrain three prominent Mamba-style restoration networks [20, 63, 96] on the all-weather setting. These experiments were conducted carefully following the settings used by other recent methods like ConvIR [13], FSNet [92], and Histoformer [65], while also respecting their original training configurations. The results are presented in Table˜14. Our MODEM achieves the best performance.
Table 14: Comparison with Mamba-like Methods.
Methods Snow100K-L Snow100K-S Outdoor RainDrop
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
MambaIR [20] 28.59 0.8729 33.34 0.9330 24.73 0.8808 32.16 0.9393
FreqMamba [96] 27.09 0.8624 33.52 0.9331 19.89 0.7519 30.60 0.9198
VmambaIR [63] 31.07 0.9168 36.35 0.9605 24.23 0.8558 32.18 0.9392
MODEM 32.52 0.9292 38.08 0.9673 33.10 0.9410 33.01 0.9434
A.8 More Visual Comparisons
Further visual comparisons are presented in Figs.˜11, 12, 13 and 14.
@@ A.8 More Visual Comparisons
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: In the abstract, we summarize our main claims and the scope of our research. In the introduction, we clearly state our contributions.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We have discussed our limitations in Section˜A.5.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper provides detailed descriptions of the proposed MODEM architecture in Section˜4, training settings, and datasets used for training and evaluation, along with the evaluation metrics in Sections˜5, A.1 and A.2.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The code will be made open-source.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper specifies key training and test details, including the datasets utilized, data splits for training/validation/testing where applicable, and the primary hyperparameters for training the MODEM framework (e.g., learning rate, batch size, optimizer type) in Sections˜5, A.5 and A.2.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: Consistent with common practice in this specific field of image restoration, error bars or detailed statistical significance tests are not explicitly reported for the main quantitative results. We follow established evaluation protocols and reporting standards used by prior state-of-the-art methods to ensure fair and direct comparisons on benchmark datasets.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Section 5 of the paper provides detailed information regarding the computer resources.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted and presented in this paper conforms, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the potential positive societal impacts of our work in both the Introduction and Conclusion sections. The nature of this research, focused on improving image quality under adverse weather, does not inherently present negative societal impacts.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The assets used in this work (e.g., code, data, models) are properly credited, and the corresponding licenses and terms of use are respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this paper does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.